OpenStudy (anonymous):

5.The members of an organization are 10 men and 11 women. What is the probability that a committee of three selected at random is all female? a-1/7 b-33/266 c-110/1,029 i got the first answer is that correct

OpenStudy (anonymous):

hmm i doubt it

OpenStudy (anonymous):

there are a couple of ways to do this one is to find the number of ways to choose 3 out of a set of 21 for your denominator, i.e. compute \[\binom{21}{3}\]

OpenStudy (anonymous):

then for them all to be female, your numerator is \[\binom{11}{3}\]

OpenStudy (anonymous):

the other way is to just count probabilty first selected is female is \[\frac{11}{21}\] then the probability that the second one is given that the first one is is \[\frac{10}{20}\]and finally the probability that the third one is given the first two are is \[\frac{9}{19}\]your answer is the product of those three numbers

OpenStudy (anonymous):

i think i got it 33/266

OpenStudy (anonymous):

i was right thanks